OpenStudy (anonymous):

Evaluate 2x-4y for x = 2 and y = 4.

OpenStudy (anonymous):

-12

OpenStudy (anonymous):

2(2)-4(4)

OpenStudy (pottersheep):

2 x 2 - 4 x 4 = 4 - 16 = -12

OpenStudy (anonymous):

-12

OpenStudy (anonymous):

2*2=4 4*4=16 4-16=-12